OpenStudy (anonymous):

what are the coordinates to the intersection between X+Y=58000 and Y=X+2000 on a graph?

OpenStudy (anonymous):

Intersection is at: (28000, 30000)

OpenStudy (anonymous):

some for why in both and set them equal to each other then solve for x. y = 58000 - x = x + 2000 ===> x = 28000, plug x back into either of the original equations y = 58000-28000 = 30000 (28000, 30000)

OpenStudy (anonymous):

LOL solve for Y in both*